how to express compressibility factor in terms of critical  pressure and critical volume

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

 

Z=\frac{p \tilde{V}}{R T}

 

where, p is the pressure, \tilde{V} is the molar volume of the gas, T is the temperature, and R is the gas constant.

For an ideal gas the compressibility factor is Z = 1 per definition

The reduced temperature and pressure are defined as:

  T_r = \frac{T}{T_c}  and  P_r = \frac{P}{P_c}

Tc and Pc are known as the critical temperature and critical pressure of a gas. They are characteristics of each specific gas with Tc being the temperature above which it is not possible to liquify a given gas and Pc is the minimum pressure required to liquify a given gas at its critical temperature.

 T_r = \frac{T}{T_c + 8} and  P_r = \frac{P}{P_c + 8}

 

now you can eliminate T and P and put in terms of critical temp and pressure.

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..


vikas askiitian expert
13 years ago

Pc = a/27b2 , Tc = 8a/27Rb , Vc = 3b

a,b are vanderwalls equation constants

Z = PV/RT = 3/8